# On the distance problem over finite p-adic rings

### v2

@@ 1 Introduction
Let $p$ be a prime, $r$ be a positive integer, and $\mathbb{Z}/p^{r}\mathbb{Z}$ be a finite $p$-adic ring. For a postive integer $n$, the “distance” between two points $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, denoted by $||\mathbf{x}-\mathbf{y}||$, is defined by
$||\mathbf{x}-\mathbf{y}||=(x_{1}-y_{1})^{2}+\cdots+(x_{n}-y_{n})^{2}\quad(\textsf{mod }p^{r}).$
Given subsets $E_{1},E_{2}\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, the distance set determined by points in $E_{1}\times E_{2}$ is denoted by $\Delta_{n,r}(E_{1},E_{2})$, i.e.,
$\Delta_{n,r}(E_{1},E_{2})=\{\|\mathbf{x}-\mathbf{y}\|:\,\mathbf{x}\in E_{1},\,\mathbf{y}\in E_{2}\}.$
@@
In the setting of finite $p$-adic rings, the Erdős-Falconer distance problem is stated as follows.
@@
This problem was initially studied in the case $r=1$, i.e. over finite fields, due to Iosevich and Rudnev [7]. More precisely, they proved that if $|E|\geq Cp^{\frac{n+1}{2}}$ for some sufficient large constant $C$, then the distance set $\Delta_{n,1}(E)$ covers the whole field. Hart, Iosevich, Koh, and Rudnev [4] indicated that the exponent $\frac{n+1}{2}$ is sharp in odd dimensions. In even dimensions, it is conjectured that the right exponent should be $n/2$. In two dimensions, Chapman, Erdogan, Hart, Iosevich, and Koh [2] proved the exponent $4/3$ by using an extension theorem associated to circles in the plane. This result was recently improved to $5/4$ by Murphy, Petridis, the first listed author, Rudnev, and Stevens in [20] by using algebraic methods and results from incidence geometry.
@@
In this paper, the first purpose is to study the distance problem in a general setting when the distance function is replaced by a much more general polynomial. In the case of the usual distance function, one of the novelties of this paper is to present a much simpler and flexible approach than that of Lichtin. Our results are essentially sharp in odd dimensions. The second purpose of this paper is to clarify the conjecture in even dimensions and provide examples to support it. In two dimensions, by developing new restriction type estimates associated to circles and orbits, with a group theoretic argument, we will prove the $4/3$-parallel result. Surprisingly, in comparison with the finite field case, in this setting, we are able to provide a large family of sets such that the distance conjecture holds. The third purpose is to study the existence/distribution of geometric/graph configurations. To be precise, we are interested in the following questions.
@@
This paper addresses the graphs of being cycles, chains, and trees, and a geometric configuration of rectangles. In contrast to initial methods/results over finite fields, the main challenge one has to deal with in this setting is to find an effective approach to obtain a uniform density which does not depend on $r$. Lichtin’s method in [15] is based on a combination of $p$-adic analysis and estimates for a class of exponential sums mod $p$. To find cancellations in exponential sums, he investigated paths of points (descending family of neighbourhoods), and his arguments are heavily relied on computing Hessian matrix of all levels. The approach we introduce in this paper is much simpler, which involves polynomial congruences and Fourier analysis in finite rings. Our arguments only rely on the first order Jacobian matrix, and avoid the second order Hessian matrix.
@@ 1.1 Results on the generalized distance sets
For a polynomial $F(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ in $n$ variables. We denote $F_{i}(\mathbf{x})=\frac{\partial F}{\partial x_{i}}(\mathbf{x})$ for all $1\leq i\leq n$, and denote $(\nabla F)(\mathbf{x})=(F_{1}(\mathbf{x}),F_{2}(\mathbf{x}),\ldots,F_{n}(\mathbf{x}))$. Our first result reads as follows.
@@
#$\#\big{\{}\mathbf{x}(\textsf{mod }p):\,$ $F(\mathbf{x})\equiv j\,(\textsf{mod }p),\,\exists~{}1\leq t\leq n,\,\,s.t.,F_{t}(\mathbf{x})\not\equiv 0\,(\textsf{mod }p),$
@@
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$, the conclusion of Theorem 1.4 holds for all integers $r,j$ with $r\geq 1$ and $(j,p)=1$, and with the constant $C$ depending only on $F$.
@@
for $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ and $s\not\equiv 0\,(\textsf{mod }p)$. And they proved that $|\Delta_{n,1}(E_{1},E_{2})|\gg p$ whenever $E_{1},E_{2}$ are subsets of $\mathbb{F}_{p}$ satisfying
@@
Merely using the approach in [11] can not lead to a result that is uniform in $r$. With the method developed in this paper, we are able to provide a generalization of the above result in the setting of finite $p$-adic rings.
@@ 1.2 The conjecture in even dimensions
In even dimensions, one might ask about the conjectured densities to guarantee that $U_{r}\subset\Delta_{n,r}(E)$ or $|\Delta_{n,r}(E)|\gg p^{r}$. We propose the following conjecture.
@@
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$ even, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$, if the density $\delta_{E_{1},E_{2}}\gg p^{-\frac{n}{2}}$, then we have $|\Delta_{n,r}(E_{1},E_{2})|\gg p^{r}$.
@@
Let $r\geq 1$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-\frac{2}{3}}$. Then $|\Delta_{2,r}(E)|\gg p^{r}$. Here the implied constant is independent of $r$.
Compared to Theorem 1.4, the exponent $2r-\frac{2}{3}$ gives a smaller threshold density, namely, of $p^{-\frac{2}{3}}$. Regarding the proof of Theorem 1.8, it is challenging if we just want to follow the methods in [2] or in [5]. On the one hand, in a recent exposition, Liao showed that the method in [5] implies too many degenerate cases, even with $r=2$, which are very hard to deal with. On the other hand, a direct computation shows that the same happens with the approach developed in [2], namely, the main difficulty arises when finding the explicit form of the sum $\sum_{j\in\mathbb{Z}/p^{r}\mathbb{Z}}\widehat{1_{C_{r,j}}}(\mathbf{m})\widehat{1_{C_{r,j}}}(\mathbf{m}^{\prime})$ for $\mathbf{m}$ and $\mathbf{m}^{\prime}$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, where $C_{r,j}$ is the circle centered at the origin of radius $j$.
In this paper, to prove Theorem 1.8, we first use a group theoretic argument, which we learned from [19, Appendix], to reduce the theorem to an extension type question for circles, then the rest is devoted to study such type estimates in the p-adic setting. The extension theorems are of independent interest and are expected to have many other applications.
@@
Let $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
@@
Let $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
@@ 1.3 Results on geometric/graph configurations
Let $F(\mathbf{x})$ be a given polynomial in $\mathbb{Z}[\mathbf{x}]$ in $n\geq 2$ variables. We call $F(\mathbf{x})$ good if it satisfies the conditions of Theorem 1.4. In the following, we obtain a number of extensions of the distance result in the setting of geometric/graph configurations in which the polynomials $F(\mathbf{x})$ are assumed to be good.
@@
Let $F$ be a good polynomial. For given $10<\delta^{\prime}<\delta<$ 1 and an integer $j$ with $(j,p)=1$, there exists $\epsilon>0$ such that the following holds. If $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2n}$, with $p^{-\frac{2n-1}{2}}\leq\epsilon$, satisfies $|E|\geq\delta p^{2rn}$, then $E$ contains at least $\delta^{\prime}p^{4rn-2}$ tuples
@@
that form rectangles of side-length $j$, i.e. $F(\mathbf{u}_{1}-\mathbf{u}_{2})=F(\mathbf{v}_{1}-\mathbf{v}_{2})\equiv j\,(\textsf{mod }p)$.
If we only want to count the number of quadruples $(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4})$ in $E$ that form a cycle, i.e., $F(\mathbf{u}_{1}-\mathbf{u}_{2})=F(\mathbf{u}_{2}-\mathbf{u}_{3})=F(\mathbf{u}_{3}-\mathbf{u}_{4})=F(\mathbf{u}_{4}-\mathbf{u}_{1})\equiv j\,(\textsf{mod }p)$, then a weaker condition is sufficient.
@@
Let $F$ be a good polynomial. There is some $C>0$ such that, if $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains cycles of length $4$ with distinct vertices and of side-length $j$.
The next result is on the existence of $k$-chains, i.e., a graph of $k+1$ vertices $\mathbf{u}_{1},\ldots,\mathbf{u}_{k+1}$ such that $F(\mathbf{u}_{i}-\mathbf{u}_{i+1})\equiv j\,(\textsf{mod }p)$ for all $11\leq i\leq k-$ 1.
@@
Let $F$ be a good polynomial, and $k\geq 1$ be an integer. There is some $C>0$ such that, if $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains $k$-chains of side-length $j$.
@@
Let $F$ be a good polynomial. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, $k\geq 1$ be an integer, and $(T,\mathbf{v})$ be a given tree with $k+1$ vertices and $k$ edges. If $\delta_{E}\gg_{k}p^{-\frac{n-1}{2}}$, then there exists $\mathbf{x}\in E$ such that the number of distinct trees $(T^{\prime},\mathbf{x})$ with vertices in $E$ and isotropic to $(T,\mathbf{v})$ is $\gg p^{rk}$.
In the rest of the paper, in some places, we may denote $q=p^{r}$ for simplicity if it causes no harm to the argument.
@@ 2 Notations and Lemmas
In this paper, the letter $p$ always denote a given prime greater than $2$. The letter $n$ denotes the dimension of the space under consideration, and $r$ is always a positive integer. The letters $m,k,l,i,j$ always denote integers. The imaginary unit is denoted by $\mathsf{i}=\sqrt{-1}$. The cardinality of a finite set $S$ is denoted by either $|S|$ or #$\#S$.
@@
$\widehat{f}(\mathbf{m})=\frac{1}{p^{rn}}\sum\limits_{\mathbf{m}\,(\textsf{mod }p^{r})}f(\mathbf{x})e_{p^{r}}(-\mathbf{m}\cdot\mathbf{x}),\quad\big{(}\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\big{)},$
@@
The following is one version of Hensel’s lemma that we will regularly applied in this paper.
@@
for any $i=1,2,\ldots,m$. Then $G_{i}(\mathbf{y}+p^{l}\mathbf{z})\equiv 0\,(\textsf{mod }p^{l+1})$ $(i=1,2,\ldots,m)$ if and only if $b_{i}+(\nabla G_{i})(\mathbf{y})\cdot\mathbf{z}\equiv 0\,(\textsf{mod }p)$ $(i=1,2,\ldots,m)$, if and only if $J_{\mathbf{G}}(\mathbf{y})\mathbf{z}\equiv-\mathbf{b}\,(\textsf{mod }p)$. The number of solutions to such a system of linear equations does not exceed $p^{n-R}$. Thus,
@@
Lemma 2.2 (Weil’s theorem).

### v3

@@ 1 Introduction
Let $p$ be a prime, $r$ be a positive integer, and $\mathbb{Z}/p^{r}\mathbb{Z}$ be a finite $p$-adic ring. For a postive integer $n$, the distance function between two points $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, denoted by $||\mathbf{x}-\mathbf{y}||:=F(\mathbf{x},\mathbf{y})$, is given by
$F(\mathbf{x},\mathbf{y})=(x_{1}-y_{1})^{2}+\cdots+(x_{n}-y_{n})^{2}\quad(\textsf{mod }p^{r}).$
Given subsets $E_{1},E_{2}\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, the distance set determined by points in $E_{1}\times E_{2}$ is denoted by $\Delta_{n,r}(E_{1},E_{2})$, i.e.
$\color[rgb]{0,0,0}\Delta_{n,r}(E_{1},E_{2})=\{F(\mathbf{x},\mathbf{y}):\,\mathbf{x}\in E_{1},\,\mathbf{y}\in E_{2}\}.$
@@
In the setting of finite $p$-adic rings, the Erdős–Falconer distance problem is stated as follows.
@@
This problem was initially studied with the distance function in the case $r=1$, i.e. over finite fields, due to Iosevich and Rudnev [7]. More precisely, they proved that if $|E|\geq Cp^{\frac{n+1}{2}}$ for some sufficient large constant $C$, then the distance set $\Delta_{n,1}(E)$ covers the whole field. Hart, Iosevich, Koh, and Rudnev [4] indicated that the exponent $\frac{n+1}{2}$ is sharp in odd dimensions. In even dimensions, it is conjectured that the right exponent should be $n/2$. In two dimensions, Chapman, Erdogan, Hart, Iosevich, and Koh [2] proved the exponent $4/3$ by using an extension theorem associated to circles in the plane. This result was recently improved to $5/4$ by Murphy, Petridis, the first listed author, Rudnev, and Stevens in [21] by using algebraic methods and results from incidence geometry.
@@
In this paper, the first purpose is to study the distance problem in a general setting when the distance function is replaced by a more general polynomial. In the case of the usual distance function, one of the novelties of this paper is to present a simpler and flexible approach than that of Lichtin. Our results are essentially sharp in odd dimensions. The second purpose of this paper is to clarify the conjecture in even dimensions and provide examples to support it. In two dimensions, by developing new restriction type estimates associated to circles and orbits, with a group theoretic argument, we will prove the $4/3$-parallel result. Surprisingly, in comparison with the finite field case, in this setting, we are able to provide a large family of sets such that the distance conjecture holds. The third purpose is to study the existence/distribution of geometric/graph configurations. To be precise, we are interested in the following configuration-type questions.
@@
This paper addresses the graphs of being cycles, chains, and trees, and a geometric configuration of rectangles. In contrast to initial methods/results over finite fields, the main challenge one has to deal with in this setting is to find an effective approach to obtain a uniform density which does not depend on $r$. Lichtin’s method in [16] is based on a combination of $p$-adic analysis and estimates for a class of exponential sums mod $p$. To find cancellations in exponential sums, he investigated paths of points (descending family of neighbourhoods), and his arguments are heavily relied on computing Hessian matrix of all levels. The approach we introduce in this paper is simpler, which involves polynomial congruences and Fourier analysis in finite rings. Our arguments only rely on the first order Jacobian matrix, and avoid the second order Hessian matrix.
@@ 1.1 Results on the generalized distance sets
For a polynomial $F(\mathbf{x})\in\mathbb{Z}[\mathbf{x}]$ in $n$ variables, we denote $F_{i}(\mathbf{x})=\frac{\partial F}{\partial x_{i}}(\mathbf{x})$ for all $1\leq i\leq n$, and denote $(\nabla F)(\mathbf{x})=(F_{1}(\mathbf{x}),F_{2}(\mathbf{x}),\ldots,F_{n}(\mathbf{x}))$. Our first result reads as follows.
@@
#$\#\big{\{}\mathbf{x}(\textsf{mod }p):\,$ $F(\mathbf{x})\equiv j\,(\textsf{mod }p),\,\exists~{}1\leq t\leq n,\,\,s.t.~{}F_{t}(\mathbf{x})\not\equiv 0\,(\textsf{mod }p),$
@@
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$ and the density threshold at least $Cp^{-\frac{n-1}{2}}$, the conclusion of Theorem 1.4 holds for all integers $r,j$ with $r\geq 1$ and $(j,p)=1$, and with the constant $C$ depending only on $F$.
@@
for $\mathbf{m}\not\equiv\mathbf{0}\,(\textsf{mod }p)$ and $s\not\equiv 0\,(\textsf{mod }p)$. And they proved that $|\Delta_{n,1}(E_{1},E_{2})|\gg p$ whenever $E_{1},E_{2}$ are subsets of $\mathbb{F}_{p}^{n}$ satisfying
@@
We observe that the approach in [11] will not lead to result which is uniform in $r$. With the method developed in this paper, we are able to provide a generalization of the above result in the setting of finite $p$-adic rings.
@@ 1.2 The conjecture in even dimensions
In even dimensions, one might ask about the conjectured densities to guarantee that $U_{r}\subset\Delta_{n,r}(E)$ or $|\Delta_{n,r}(E)|\gg p^{r}$. In this paper, we propose the following.
@@
Let $F(\mathbf{x})=\sum\nolimits_{i=1}^{n}a_{i}x_{i}^{k}$ be a polynomial in $\mathbb{Z}[\mathbf{x}]$ with $n\geq 2$ even, $k\geq 2$, and $a_{i}\neq 0$ for all $1\leq i\leq n$. Then, for any sufficiently large prime $p$, if $r$ is a positive integer and $E_{1},E_{2}$ are subsets of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with the density $\delta_{E_{1},E_{2}}\gg p^{-\frac{n}{2}}$, we have $|\Delta_{n,r}(E_{1},E_{2})|\gg p^{r}$.
@@
Let $F$ be the distance function, $p$ be an odd prime, and $r\geq 1$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-\frac{2}{3}}$. Then $|\Delta_{2,r}(E)|\gg p^{r}$. Here the implied constant is independent of $r$.
Compared to Theorem 1.4, we have a smaller threshold density, namely, of $p^{-\frac{2}{3}}$. Regarding the proof of Theorem 1.8, it is challenging if we just want to follow the methods in [2] or in [5]. On the one hand, in a recent exposition, Liao showed that the method in [5] implies too many degenerate cases, even with $r=2$, which are very hard to deal with. On the other hand, a direct computation shows that the same happens with the approach developed in [2], namely, the main difficulty arises when finding the explicit form of the sum $\sum_{j\in\mathbb{Z}/p^{r}\mathbb{Z}}\widehat{1_{C_{r,j}}}(\mathbf{m})\widehat{1_{C_{r,j}}}(\mathbf{m}^{\prime})$ for $\mathbf{m}$ and $\mathbf{m}^{\prime}$ in $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, where $C_{r,j}$ is the circle centered at the origin of radius $j$.
In this paper, to prove Theorem 1.8, we first use a group theoretic argument, which we learned from [20, Appendix], to reduce the theorem to an extension type question for circles, then the rest is devoted to study such type estimate in the p-adic setting. The extension theorems, we obtain in Section 4, are of independent interest and are expected to have further applications.
@@
Let $F$ be the distance function, $p$ be an odd prime, and $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
@@
Let $F$ be the distance function, $p$ be an odd prime, and $r\geq 2$ be an integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2}$ with $\delta_{E}\gg p^{-1}$. Assume that
@@ 1.3 Results on geometric/graph configurations
Let $F(\mathbf{x})$ be a given polynomial in $\mathbb{Z}[\mathbf{x}]$ in $n\geq 2$ variables. We call $F(\mathbf{x})$ good if it satisfies the conditions of Theorem 1.4. In the following, we prove several extensions of the distance result in the setting of geometric/graph configurations in which the polynomials $F(\mathbf{x})$ are assumed to be good.
@@
Let $F$ be a good polynomial and $10<\delta^{\prime}<\delta<$ 1 be given numbers. There exists $\epsilon>0$ such that the following holds. For any odd prime $p$, any positive integer $r$, and any integer $j$ with $(j,p)=1$, if $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{2n}$, with $p^{-\frac{2n-1}{2}}\leq\epsilon$, satisfies $\delta_{E}\geq\delta$, then $E$ contains at least $\delta^{\prime}p^{4rn-2}$ tuples
@@
that form rectangles of side-length $j$, i.e. $F(\mathbf{u}_{1}-\mathbf{u}_{2})\equiv F(\mathbf{v}_{1}-\mathbf{v}_{2})\equiv j\,(\textsf{mod }p^{r})$.
If we only want to count the number of quadruples $(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4})$ in $E$ that form a cycle, i.e., $F(\mathbf{u}_{1}-\mathbf{u}_{2})=F(\mathbf{u}_{2}-\mathbf{u}_{3})=F(\mathbf{u}_{3}-\mathbf{u}_{4})=F(\mathbf{u}_{4}-\mathbf{u}_{1})\equiv j\,(\textsf{mod }p^{r})$, then a weaker condition is sufficient.
@@
Let $F$ be a good polynomial and $p$ be an odd prime. There is some $C>0$ such that, if $r$ is a positive integer, $j$ is an integer with $(j,p)=1$, and $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains cycles of length $4$ with distinct vertices and of side-length $j$.
The next result is on the existence of $k$-chains, i.e. a graph of $k+1$ vertices $\mathbf{u}_{1},\ldots,\mathbf{u}_{k+1}$ such that $F(\mathbf{u}_{i}-\mathbf{u}_{i+1})\equiv j\,(\textsf{mod }p^{r})$ for all $11\leq i\leq k-$ 1.
@@
Let $F$ be a good polynomial, $p$ be an odd prime, and $k\geq 1$ be an integer. There is some $C>0$ such that, if $r$ is a positive integer, $j$ is an integer with $(j,p)=1$, and $E$ is a subset of $(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$ with $\delta_{E}\geq Cp^{-\frac{n-1}{2}}$, then $E$ contains $k$-chains of side-length $j$.
@@
Let $F$ be a good polynomial, $p$ be an odd prime, and $r$ be a positive integer. Let $E\subset(\mathbb{Z}/p^{r}\mathbb{Z})^{n}$, $k\geq 1$ be an integer, and $(T,\mathbf{v})$ be a given tree with $k+1$ vertices and $k$ edges. If $\delta_{E}\gg_{k}p^{-\frac{n-1}{2}}$, then there exists $\mathbf{x}\in E$ such that the number of distinct trees $(T^{\prime},\mathbf{x})$ with vertices in $E$ and isotropic to $(T,\mathbf{v})$ is $\gg p^{rk}$.
@@ 2 Notations and Lemmas
In this paper, the letter $p$ always denote a given prime greater than $2$. The letter $n$ denotes the dimension of the space under consideration, and $r$ is always a positive integer. The letters $m,k,l,i,j$ always denote integers. The imaginary unit is denoted by $\mathsf{i}=\sqrt{-1}$. The cardinality of a finite set $S$ is denoted by either $|S|$ or #$\#S$. By $X\ll Y$, it means there exists some positive constant $C$ such that $X\leq CY$. If $C=C(k)$, for some parameter $k$, then we write $X\ll_{k}Y$.
@@
$\widehat{f}(\mathbf{m})=\frac{1}{p^{rn}}\sum\limits_{\mathbf{x}\,(\textsf{mod }p^{r})}f(\mathbf{x})e_{p^{r}}(-\mathbf{m}\cdot\mathbf{x}),\quad\big{(}\mathbf{m}\in(\mathbb{Z}/p^{r}\mathbb{Z})^{n}\big{)},$
@@
The next lemma is one version of Hensel’s lemma that will be regularly applied in this paper.
@@
for any $i=1,2,\ldots,m$. Then $G_{i}(\mathbf{y}+p^{l}\mathbf{z})\equiv 0\,(\textsf{mod }p^{l+1})$ $(i=1,2,\ldots,m)$ if and only if $b_{i}+(\nabla G_{i})(\mathbf{y})\cdot\mathbf{z}\equiv 0\,(\textsf{mod }p)$ $(i=1,2,\ldots,m)$, if and only if $J(\mathbf{G})|_{\mathbf{y}}\mathbf{z}\equiv-\mathbf{b}\,(\textsf{mod }p)$. The number of solutions to such a system of linear equations does not exceed $p^{n-R}$. Thus,
@@
In the following, we collect and prove some results involving exponential sums or cardinarlity of varieties.
Lemma 2.2 (Weil’s theorem, Theorem 5.38, [12]).
@@
Here $X=o_{F}(Y)$, it means $X/Y\rightarrow 0$ with the convergence rate depending on $F$.